Assume the following for the town of Boone: it has a total population of 40,000 people, of which 4,000 are under 16 years of age or are institutionalized; 5,000 are full-time students who are not employed and are not seeking work; and 30,000 are employed. The rest of the people are out of work but have been actively seeking work within the past four weeks.

What is Boone's unemployment rate?
Answer #1

Labor force = total population - under age - student not seeking work = 40000 - 4000 - 5000 = 31000

Employed = 30000

Unemployed = labor force - employed = 31000 - 30000 = 1000

unemployment rate = 1000 / 31000 * 100 = 3.226% = 3.23%